OpenStudy (anonymous):

Identify the vertical asymptotes of f(x) = x-4/x^2+13x+36

OpenStudy (anonymous):

Parenthesis please... I'll rewrite it for what I think you meant

OpenStudy (anonymous):

\[(x-4)/(x^2+13x+36)=\frac{x-4}{x^2+13x+36}\] This?

OpenStudy (anonymous):

yes thats it

OpenStudy (anonymous):

What you wrote initially was: \[x-\frac{4}{x^2}+13x+36\] So yeah, definitely NOT the same thing :-D It's kinda important, just for future reference. Here is your function factored: \[ \frac{(x-4)}{(x+4)(x+9)} \] Check: x*x=x^2 4x+9x=13x 4*9=36 Vertical asymptotes are x = some constant, numerical # You want to solve the denominator for when it's equal to zero (x+4)(x+9)=0 See if you can find the answer now :-)

OpenStudy (anonymous):

If you graph it on a TI-84 or something like that, when you trace along the function to those x values you'll get errors or blanks for the "y=" display.